# template.bib

### v1

@@ 5 Discussion
While SHAP provides detailed, feature-level explanations, its effectiveness in capturing the true importance of features may be impacted by feature dependencies [aasExplainingIndividualPredictions2021]. SHAP’s default independence assumption may lead to diluted feature importance in scenarios with correlated features, such as the kinematic chain in human motion. GradCAM, with its spatial focus and faster runtime, is more suited to applications where visual feedback and quick assessments are needed. The two methods can complement each other, with SHAP giving the ”why“ behind model decisions and Grad-CAM giving the ”where.“ In conclusion, the explainability of an ML model should always be tailored to different stakeholders, providing each with an appropriate level of detail based on their specific needs [koricaExplainableArtificialIntelligence2021].
@@ 6 Conclusion
This paper systematically compared SHAP and GradCAM within the context of GCNs and HAR using skeleton data from two real-world datasets. Our findings highlight that these two explainability methods differ in their approach and output, each offering different benefits depending on the specific requirements of the individual use case.
SHAP provides a detailed, feature-level explanation, making it practical in scenarios where it is crucial to understand each individual body’s key points’ contribution in combination with the input features. It offers a more profound, finer view of how specific features impact the model’s decisions, which is essential for trust and accountability in sensitive domains. This may be required in healthcare applications such as CP prediction, where decisions must be transparent and explainable to clinicians. However, its computational intensity and longer runtime make it less suitable for real-time applications.
In contrast, GradCAM provides quicker, spatially oriented explanations, showing which regions of the body are most influential in the model’s decision at a specific time frame. Its speed and simplicity make it a better choice for real-time assessments or scenarios where quick and more straightforward visual feedback is necessary. However, its lack of granularity and inability to capture complex feature interactions from earlier layers limits its effectiveness in circumstances demanding deeper insights.
Nevertheless, our study demonstrates that SHAP and GradCAM should be viewed as complementary tools rather than competing methods. SHAP provides the ”why“ behind decisions, while GradCAM offers the ”where.“ Depending on the specific demands of the task, whether it is the need for quick, spatial understanding or a thorough, feature-level interpretation, these methods can be used in parallel to provide a more complete and interpretable explanation of the model behavior on the dataset. For future work, hybrid approaches that combine the strengths of both SHAP and GradCAM should be explored. We can enhance these systems’ safety, transparency, and trustworthiness by tailoring explainability techniques to specific HAR models and datasets, particularly in high-risk domains like healthcare.

### v2

@@ 5 Discussion
While SHAP provides detailed, feature-level explanations, its effectiveness in capturing the true importance of features may be impacted by feature dependencies [aasExplainingIndividualPredictions2021]. SHAP’s default independence assumption may lead to diluted feature importance in scenarios with correlated features, such as the kinematic chain in human motion. Grad-CAM, with its spatial focus and faster runtime, is more suited to applications where visual feedback and quick assessments are needed. The two methods can complement each other, with SHAP giving the “why” behind model decisions and Grad-CAM giving the “where”. In conclusion, the explainability of an ML model should always be tailored to different stakeholders, providing each with an appropriate level of detail based on their specific needs [koricaExplainableArtificialIntelligence2021].
6 Limitations and Future Work
While our approach of comparing Grad-CAM with SHAP shows the individual strengths and weaknesses, limitations exist, which open possibilities for future work.
Our comparison study is restricted to the two explanation methods, SHAP and Grad-CAM, which are the typical approaches applied for GCNs within the HAR domain. While these two techniques are the common choice when explaining skeleton data, other explanation methods should be discovered. This can include the predecessor of Grad-CAM, Grad-CAM++ [chattopadhyayGradCAMImprovedVisual2018], a generalized approach that improves localization capability. In particular, Grad-CAM++ may produce more localized explanations, which could be validated with the perturbation experiment. Another possible explanation method that could be used is LIME [ribeiroWhyShouldTrust2016]. However, perturbing skeleton data is challenging since naive alterations of the sequence can produce unrealistic or biomechanically implausible actions and movements. One potential solution is to define kinematic sound perturbations that respect natural joint movement constraints by simulating only plausible variations or applying domain-specific transformations. This highlights a key issue in applying LIME to skeleton-based HAR: ensuring that the perturbed data remains realistic so that the local surrogate model returns valid insights. Future work should, therefore, incorporate the mentioned explanation methods to provide the bigger picture beyond the two applied XAI methods within this work.
In our current work, we apply the two quantitative metrics, PGI and PGU, to assess the reliability of the obtained explanations [agarwalOpenXAITransparentEvaluation2024], acknowledging that there exists more metrics to assess the reliability of an explanation. The PGI and PGU metrics are particularly useful in measuring the faithfulness of the explanation when no ground truth is available, which is the case within our approach. However, future work should explore metrics to measure the explanation stability, such as Relative Input Stability (RIS), Relative Representation Stability (RRS), and Relative Output Stability (ROS), to measure the explanation change when perturbing parts of the model parameters [agarwalRethinkingStabilityAttributionbased2022]. With these additional metrics, the stability of the explanation can be assessed, contributing to a multifaceted view of the reliability of the obtained explanations.
Another promising avenue for future research is the incorporation of human-centric explanations, which are interrogative. In [sporsemCliniciansDontKnow2025], the authors show that clinicians prefer such interrogative explanations, where they can explore the output of the explanations on their own. By providing multiple explanation methods, clinicians can choose the most suited one for their specific use case. Therefore, it is essential to communicate the individual distinctions of each explanation method, such as the spatial orientation of Grad-CAM or the granular explanation capabilities of SHAP, for which this comparative study lays the groundwork.
Depending on the specific demands of the task, whether it is the need for quick, spatial understanding or a thorough, feature-level explanation, these methods can be used in parallel to provide a complete explanation of the model behavior on the dataset. For future work, hybrid approaches that combine the strengths of both SHAP and Grad-CAM should be explored. This could be achieved through a combined XAI score leveraging the Grad-CAM for spatial explanation and SHAP for feature-level explanation, resulting in an overall score that balances spatial and feature-level explanations.
@@ 7 Conclusion
This paper systematically compared SHAP and Grad-CAM within the context of GCNs and HAR using skeleton data from two real-world datasets. Our findings highlight that these two explainability methods differ in their approach and output, each offering different benefits depending on the specific requirements of the individual use case.
SHAP provides a detailed, feature-level explanation, making it practical in scenarios where it is crucial to understand each individual body’s key points’ contribution in combination with the input features. It offers a more profound, finer view of how specific features impact the model’s decisions, which is essential for granular insights in sensitive domains. This may be required in healthcare applications such as CP prediction, where model decisions must be transparent and explainable to clinicians. However, its computational intensity and longer runtime make it less suitable for real-time applications.
In contrast, Grad-CAM provides quicker, spatially oriented explanations, showing which regions of the body are most influential in the model’s decision at a specific time frame. Its speed and simplicity make it a better choice for real-time assessments or scenarios where quick and more straightforward visual feedback is necessary. However, its lack of granularity and inability to capture complex feature interactions from earlier layers limits its effectiveness in circumstances demanding granular feature-level insights.
Finally, our study demonstrates that SHAP and Grad-CAM should be viewed as complementary tools rather than competing methods. SHAP provides the “why” behind decisions, while Grad-CAM gives the “where”.